a partcle projected from ground moves at angle 45 degree one second after projection and speed is minimum two seconds after projection. The angle of projection of particle is (neglect the air resistance) :
  1. tan-1 (3)
  2. tan-1 (2)
  3. tan-1 (root2)
  4. tan-1 (4)

Grade:11

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,

speed is minimum after 2 sec means speed=vcos theta
after 1 second
v'cos45=vcos theta
also
v'sin45-g*1=0
=>v'sin45=g
=>v'=gsqrt2
also
vsin theta-g*1=v'sin45=g
=>vsin theta=2g
=>tan theta=2g/v'cos45=2

Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty